jerry spent 2.5 h biking up mount lowe, rested at the top for 30 min, and biked down in 1.5 hr. how far did he bike if his rate of ascent was 3 km/h less than his rate of descent?
Answer by stanbon(75887)

jerry spent 2.5 h biking up mount lowe, rested at the top for 30 min, and biked down in 1.5 hr. how far did he bike if his rate of ascent was 3 km/h less than his rate of descent?
----------------
Biking up DATA:
time = 2.5 hr. ; rate= x-3 km/hr ; distance = rt= 2.5(x-3) km
-----------
Biking down Data:
time = 1.5 hr; rate = x km/hr ; distance = 1.5x km
---------
EQUATION:
1.5x=2.5x-7.5
-x=-7.5
x=7.5 km/hr (Biking down rate)
Distance= 7.5(1.5)=11.25 km (Biking down distance Biking up distance)
Cheers,
Stan H.